# Random local access for sampling $k$-SAT solutions

### v2

@@ Abstract
We present a sublinear time algorithm that gives random local access to the uniform distribution over satisfying assignments to an arbitrary $k$-SAT formula $\Phi$, at exponential clause density. Our algorithm provides memory-less query access to variable assignments, such that the output variable assignments consistently emulate a single global satisfying assignment whose law is close to the uniform distribution over satisfying assignments to $\Phi$.
Such models were formally defined (for the more general task of locally sampling from exponentially sized sample spaces) in 2017 by Biswas, Rubinfeld, and Yodpinyanee, who studied the analogous problem for the uniform distribution over proper $q$-colorings. This model extends a long line of work over multiple decades that studies sublinear time algorithms for problems in theoretical computer science. Random local access and related models have been studied for a wide variety of natural Gibbs distributions and random graphical processes. Here, we establish feasibility of random local access models for one of the most canonical such sample spaces, the set of satisfying assignments to a $k$-SAT formula.
@@ 1 Introduction
We give a more precise definition of random local access models in Section 2. Random local access models were formally introduced in work of Biswas, Rubinfeld, and Yodpinyanee [4], motivated by a long line of work over the past decades studying sublinear time algorithms for problems in theoretical computer science. The authors of [4] studied a different natural Gibbs distribution, the uniform distribution over proper $q$-colorings of a given $n$-vertex graph with maximum degree $\Delta$. By adapting previously studied classical and parallel sampling algorithms for $q$-colorings, they were able to construct a local access algorithm to a random proper $q$-coloring when $q\geq 9\Delta.$ The related problem of sampling partial information about huge random objects was pioneered much earlier in [15, 16]; further work in [27, 8] considers the implementation of different random graph models. Random local access and related (e.g., parallel computation) models have also been recently studied for several other random graphical processes and Gibbs distributions (cf. [4, 3, 25, 23]).
@@
The above model for local access to random samples extends a rich line of work studying local computation algorithms (LCAs), originally introduced in works of Rubinfield, Tamir, Vardi, and Xie [28] and Alon, Rubinfield, Vardi, and Xie [1]. A primary initial application of LCAs in [28] was locally constructing a satisfying assignment to a given $(k,d)$-formula $\Phi$ when $d\lesssim 2^{k/10}$. Similar to the desired properties of random local access, LCA implements query access to a large solution space in sublinear time using local probes. However, rather than sampling from a given distribution, LCA only provides local access to some consistent valid solution in the desired solution space that may otherwise be completely arbitrary in nature.
@@
In this work, we present a polylogarithmic time algorithm that gives random local access to the uniform distribution over satisfying assignments of an arbitrary $k$-SAT formula $\Phi$ at exponential clause density. The following is an informal version of our main result; a more formal version can be found at 5.3.
Theorem 1.1 (Main theorem: informal).
There exists a random local access algorithm $\mathcal{A}$, such that for any $(k,d)$-formula $\Phi=(V,\mathcal{C})$ that satisfies
$d\lesssim 2^{k/400}$ (1)
and any variable $v\in V$, $\mathcal{A}$ can be used to sample a single assignment $\widehat{\mu}_{v}\in\{0,1\}$ in $\mathrm{polylog}(n)$ time in expectation such that for all $c>0$,
$d_{\operatorname{\mathsf{TV}}}(\mathsf{Law}(\widehat{\mu}_{v}),\mu_{v})\leq\frac{1}{n^{c}}.$
@@
Similar to [4], the proof of 1.1 adapts some of the algorithmic tools used to study parallel and distributed sampling. The proof also builds upon the work of Moitra [24] and Feng–Guo–Yin–Zhang [10] on sampling from bounded degree $k$-CNFs in polynomial time. The authors of [24, 10] both critically took advantage of a global variable marking (see 2.9) to better condition the marginal distributions of variables. Such an approach allows for a subset of the variable assignments to be sampled with ease; the resulting shattering of the solution space conditioned on such a partial assignment then allows one to efficiently complete the random satisfying assignment. These initial approaches have been extended and strengthened to nearly linear time algorithms that succeed for larger maximum variable degree in a flurry of recent work (c.f. [24, 10, 20, 17, 13, 6, 12, 29, 5]).
@@
Recently, Feng, Guo, Wang, Wang, and Yin [9] used a recursive sampling scheme to simulate the systematic scan projected Glauber dynamics via a strategy termed coupling towards the past, which they used to derandomise several MCMC algorithms for sampling from CSPs. Additionally, recent work of He, Wang and Yin [18] used a recursive sampling scheme to sample $k$-SAT solutions. Their work immediately gives a sublinear (in fact, expected constant time) algorithm for sampling the assignment of a single variable one time; however, this work does not immediately extend out of the box to a random local access model that enjoys the desired consistency and memory-less properties if multiple variables are sampled. Recursive sampling schemes have also been recently used to make progress on analyzing and designing fast sampling algorithms for a variety of Gibbs distributions (cf. [18, 6, 2]). As noted earlier, such schemes have been particularly useful for derandomising and constructing parallel and distributed versions of many popular MCMC algorithms for sampling solutions to CSPs [9, 11, 17, 18].
@@
We use these algorithms to carefully select and sample a small number of auxiliary variable assignments on an as-needed basis, showing that bad cases can be reduced to small calculations after a sublinear number of steps. Our proof of 1.1 requires a novel adaptation of sampling techniques to avoid requiring global information or complete passes of the variable set; we show that we can perform analogous operations locally or leverage local uniformity properties to circumvent them altogether in the course of locally sampling a variable assignment. Importantly, we demonstrate how to execute local sampling in an oblivious, consistent fashion, so that the algorithm need not retain any memory and that variables need not be sampled in any particular order.
Acknowledgements
We would like to thank Kuikui Liu and Ronitt Rubinfeld for helpful discussions and feedback on a draft. We would like to thank the anonymous referee for identifying an issue with a previous version of this article. NM was supported by the Hertz Graduate Fellowship and the NSF Graduate Research Fellowship Program #2141064. This material includes work supported by the NSF Grant DMS-1928930, while the authors were in residence at SLMath during the Spring 2025 semester.
@@ 2 Preliminaries
Throughout, let $\Phi=(V,\mathcal{C})$ be a $k$-CNF on variable set $V=\{v_{1},\ldots,v_{n}\}$ with associated collection of clauses $\mathcal{C}$. We further assume that $\Phi$ is a $(k,d)$-formula; that is, each variable $v_{i}$ appears in at most $d$ clauses. For every clause $C\in\mathcal{C}$, let $\operatorname{vbl}(C)$ denote the collection of variables in the clause $C$. We further define $\Delta:=\max_{C\in\mathcal{C}}|\{C^{\prime}\in\mathcal{C}:\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset\}|$, so in particular $\Delta\leq kd$.
@@
Definition 2.2 (Dependency hypergraph).
Given a $(k,d)$-formula $\Phi$, let $H_{\Phi}=(V,\mathcal{E})$ be the dependency hypergraph (with multiple edges allowed), where $V$ is the set of variables and =$\mathcal{E}=\{\operatorname{vbl}(C):C\in\mathcal{C}\}$ is the collection of clauses of $\Phi$ viewed as $k$-sets of variables.
Definition 2.3 (Partial assignment).
@@
Let $H_{\Phi}^{\sigma}$ be the associated (not necessarily $k$-uniform) hypergraph to $\Phi^{\sigma}$. For variable $v\in V\setminus\Lambda$, let $\Phi_{v}^{\sigma}$ denote the connected component of $\Phi^{\sigma}$ to which $v$ belongs.
Definition 2.5 (Marginal distribution).
@@ 2.1 The random local access model and local computation algorithms
In this work, we construct a query-oblivious LCA for an intermediate step in our construction of the random local access model (described in more detail in Section 2.2). We thus precisely define both LCAs and random local access models below.
Definition 2.6 (Local computation algorithm).
@@
  • For a given input $\Pi$, upon being queried with $F_{i}$, the oracle with probability $1-o(1)$ uses at most $t(n)$ resources (where $t(n)$ is sublinear in $n$) to return the value $\mathcal{A}(\Pi,\mathbf{R},F_{i}(Y))$ for some specific $Y\in\Omega$.
@@ 2.2 Marking and a query-oblivious LCA
As noted in the introduction, the Lovász local lemma guarantees the existence of a satisfying assignment to any $(k,d)$-formula $\Phi$ if $d\leq(2ek)^{-1}2^{k}.$ Furthermore, the Moser-Tardos algorithm [26] gives a simple linear-time algorithm to construct such a satisfying assignment in the above $(k,d)$-regime:
[…]
@@
  • For every clause $C$, $|\operatorname{vbl}(C)\backslash\mathcal{M}|\gtrsim k$. This large non-intersection preserves the local uniformity of the marginal distributions of the as yet-unsampled variables in $\mathcal{M}$.
@@
Definition 2.9 ($\alpha$-marking).
Given $(k,d)$-formula $\Phi$ and constant $\alpha\in(0,1/2)$, we say that $\mathcal{M}\subseteq V$ is an $\alpha$-marking if for every $C\in\mathcal{C}$, $|\operatorname{vbl}(C)\cap\mathcal{M}|\geq\alpha k$ and $|\operatorname{vbl}(C)\backslash\mathcal{M}|\geq\alpha k$.
@@
The construction of $\mathsf{IsMarked}(\cdot)$ and the verification that it is a query-oblivious LCA draws inspiration from the approach in [1] to obtain an oblivious LCA for hypergraph $2$-coloring. We defer the proof of Theorem 2.10 to Appendix D.
@@ 3 A random local access algorithm for $k$-SAT solutions
In this section, we introduce the main algorithm (Algorithm 1) that locally accesses the assignment of a variable $v\in V$ in a uniformly random satisfying assignment $\sigma$ to $\Phi$.
Recall from 1.1 that given a $(k,d)$-formula $\Phi=(V,\mathcal{C})$, variable $v\in V$, and any constant $c>0$, we wish to come up with a random local access algorithm $\mathcal{A}$ such that (1) the expected running time of $\mathcal{A}$ is $\mathrm{poly}\log(n)$, and (2) the output $\widehat{\mu}_{v}$ of $\mathcal{A}$ for every $v\in V$ satisfies $d_{\operatorname{\mathsf{TV}}}(\textsf{Law}(\widehat{\mu}_{v}),\mu_{v})\leq\frac{1}{n^{c}}$. As a high level description, given the above inputs, Algorithm 1 performs the following:
@@
    Locally decides whether $v$ lies in an $\alpha$-marking of $\Phi$ using $\mathsf{IsMarked}(v)$ (Theorem 2.10), such that the responses for all $v\in V$ yield a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
@@
    After obtaining the connected component $\Phi_{v}^{\sigma}$, we call $\mathsf{UniformSample}$ (Theorem 3.1) to sample a uniformly random satisfying assignment to $\Phi_{v}^{\sigma}$ and extend $\sigma$. We then output $\sigma(v)$.
@@
Algorithm 1 is our main routine; it has several subroutines $\mathsf{IsMarked}$, $\mathsf{MarginSample}$, $\mathsf{Conn}$, and $\mathsf{UniformSample}$ that we now describe. Recall that $\mathsf{IsMarked}$ has been discussed in Theorem 2.10. We now introduce the algorithm $\mathsf{UniformSample}$ given by the work of He–Wang–Yin [18].
@@
As seen in Theorem 3.1, for an $n$-variable $(k,d)$-CNF formula, the algorithm has running time $O(n)$. However, as we will only apply $\mathsf{UniformSample}$ to connected components of size $\mathrm{poly}\log(n)$ (that are shattered by the partial assignment on the marked variables), the execution of $\mathsf{UniformSample}$ in Line 6 of Algorithm 1 will only takes polylogarithmic time.
We will define subroutines $\mathsf{MarginSample}$ and $\mathsf{Conn}$ below and show that they satisfy the desired correctness and efficiency properties, beginning by verifying $\mathsf{MarginSample}$ has the desired properties in Section 4.
@@
Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking as in 2.9. Fix positive constant $c>0$. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$, $236ed^{3}k^{4}\cdot 0.6^{\alpha k}\leq 1/$ 2, and $2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{\alpha k}}}\leq 0.9$. Then there exists a random local access algorithm $\mathsf{MarginSample}(\cdot)$ such that for every $u\in\mathcal{M}$, $\mathsf{MarginSample}(u)$ returns a random value in $\{0,1\}$ that satisfies the following:
@@
We also require $\mathsf{Conn}$ to be correct and have low expected cost, verified in Appendix C.
@@
Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking as in 2.9. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$ and $d\leq 2^{\alpha k/4}$. Then there exists a random local access algorithm $\mathsf{Conn}(\cdot)$ such that for every $v\in V\setminus\mathcal{M}$, $\mathsf{Conn}(v)$ returns the connected component in $\Phi^{\sigma}$ that contains $v$, where $\sigma$ the partial assignment given by $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$. Moreover, for every $v\in V\setminus\mathcal{M}$, the expected cost of $\mathsf{Conn}(v)$ is $\mathrm{poly}\log(n)$.
4 Proof of Theorem 3.2
In this section we show Theorem 3.2. Throughout, let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking with $|\mathcal{M}|=m$. We introduce a local access algorithm $\mathsf{MarginSample}$ that satisfies the key property that the joint distribution of outputs $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ consistently follows the marginal distribution $\mu_{\mathcal{M}}$.
In particular, $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ will simulate the output of a systematic scan Glauber dynamics on the marked variables. We begin by recalling some basics of Markov chains and the projected Glauber dynamics Markov chain for sampling $k$-SAT solutions, introduced in [10].
Definition 4.1.
Algorithm 2 Projected Glauber dynamics
      Output: a random assignment $X_{*}\in\{0,1\}^{\mathcal{M}}$
[…]
2:for $t=1,\dots,T$ do
[…]
6:Return $X_{*}=X_{T}$
We contrast Algorithm 2 with the systematic scan version of the projected Glauber dynamics. Instead of resampling a random variable at time $t$, in the systematic scan Glauber dynamics, we always resample the variable $u_{i(t)}$ at time $t$, where $i(t)$ is a deterministic function of $t$.
Definition 4.2.
Let =$\mathcal{M}=\{u_{1},\ldots,u_{m}\}$ denote the marked variables (so $|\mathcal{M}|=m$).
@@
Input: a $k$-CNF formula $\Phi=(V,\mathcal{C})$, a set of marked variables $\mathcal{M}\subseteq V$, time $T$, and an ordering =$\mathcal{M}=\{u_{1},\ldots,u_{m}\}$
@@
[…]

### v3

@@ Abstract
We present a sublinear time algorithm that gives random local access to the uniform distribution over satisfying assignments to an arbitrary $k$-SAT formula $\Phi$, at exponential clause density. Our algorithm provides memory-less query access to variable assignments, such that the output variable assignments consistently emulate a single global satisfying assignment whose law is close to the uniform distribution over satisfying assignments to $\Phi$. Random local access and related models have been studied for a wide variety of natural Gibbs distributions and random graphical processes. Here, we establish feasibility of random local access models for one of the most canonical such sample spaces, the set of satisfying assignments to a $k$-SAT formula.
Our algorithm proceeds by leveraging the local uniformity of the uniform distribution over satisfying assignments to $\Phi$. We randomly partition the variables into two subsets, so that each clause has sufficiently many variables from each set to preserve local uniformity. We then sample some variables by simulating a systematic scan Glauber dynamics backward in time, greedily constructing the necessary intermediate steps. We sample the other variables by first conducting a search for a polylogarithmic-sized local component, which we iteratively grow to identify a small subformula from which we can efficiently sample using the appropriate marginal distribution. This two-pronged approach enables us to sample individual variable assignments without constructing a full solution.
keywords:
sublinear time algorithms, random generation, $k$-SAT, local computation
category:
[…]
@@ 1 Introduction
We give a more precise definition of random local access models in Section˜2. Random local access models were formally introduced in work of Biswas, Rubinfeld, and Yodpinyanee [5], motivated by a long line of work over the past decades studying sublinear time algorithms for problems in theoretical computer science. The authors of [5] studied a different natural Gibbs distribution, the uniform distribution over proper $q$-colorings of a given $n$-vertex graph with maximum degree $\Delta$. By adapting previously studied classical and parallel sampling algorithms for $q$-colorings, they were able to construct a local access algorithm to a random proper $q$-coloring when $q\geq 9\Delta.$ The related problem of sampling partial information about huge random objects was pioneered much earlier in [15, 16]; further work in [29] considers the implementation of different random graph models. Random local access and related (e.g., parallel computation) models have also been recently studied for several other random graphical processes and Gibbs distributions (cf. [5, 4, 27, 24]).
@@
The above model for local access to random samples extends a rich line of work studying local computation algorithms (LCAs), originally introduced in works of Rubinfield, Tamir, Vardi, and Xie [30] and Alon, Rubinfield, Vardi, and Xie [1]. Local computation algorithms are widely used in parallel and distributed computing, with notable practical success in areas such as graph sparsification [3], load balancing [25], and sublinear-time coloring [1]. A primary initial application of LCAs in [30] was locally constructing a satisfying assignment to a given $(k,d)$-formula $\Phi$ when $d\lesssim 2^{k/10}$. Similar to the desired properties of random local access, LCA implements query access to a large solution space in sublinear time using local probes. However, rather than sampling from a given distribution, LCA only provides local access to some consistent valid solution in the desired solution space that may otherwise be completely arbitrary in nature.
@@
In this work, we present a polylogarithmic time algorithm that gives random local access to the uniform distribution over satisfying assignments of an arbitrary $k$-SAT formula $\Phi$ at exponential clause density (i.e., the number of occurrences of each variable is bounded by an exponential function of $k$). The following is an informal version of our main result; a more formal version can be found at ˜5.3.
[…]
There exists a random local access algorithm $\mathcal{A}$ that satisfies the following conditions. For any $(k,d)$-formula $\Phi=(V,\mathcal{C})$ with maximum degree $d\lesssim 2^{k/400}$ and any variable $v\in V$, we can use $\mathcal{A}$ to sample a single assignment for $v$, $\mathcal{A}(v)\in\{0,1\}$ in expected polylogarithmic time (in $|V|$), such that the distribution of $\mathcal{A}(v)$ is the same as the marginal distribution of $\sigma(v)$, where $\sigma$ is a uniformly random satisfying assignment to $\Phi$.
@@
Similar to [5], the proof of Section˜1 adapts some of the algorithmic tools used to study parallel and distributed sampling. The proof also builds upon the work of Moitra [26] and Feng–Guo–Yin–Zhang [10] on sampling from bounded degree $k$-CNFs in polynomial time. The authors of [26, 10] both critically took advantage of a global variable marking (see Section˜2.2) to better condition the marginal distributions of variables. Such an approach allows for a subset of the variable assignments to be sampled with ease; the resulting shattering of the solution space conditioned on such a partial assignment then allows one to efficiently complete the random satisfying assignment. These initial approaches have been extended and strengthened to nearly linear time algorithms that succeed for larger maximum variable degree in a flurry of recent work (c.f. [26, 10, 20, 17, 13, 7, 12, 32, 6]).
@@
Recently, Feng, Guo, Wang, Wang, and Yin [9] used a recursive sampling scheme to simulate the systematic scan projected Glauber dynamics via a strategy termed coupling towards the past, which they used to derandomise several Markov chain Monte Carlo (MCMC) algorithms for sampling from CSPs. Additionally, recent work of He, Wang and Yin [18] used a recursive sampling scheme to sample $k$-SAT solutions. Their work immediately gives a sublinear (in fact, expected constant time) algorithm for sampling the assignment of a single variable one time; however, this work does not immediately extend out of the box to a random local access model that enjoys the desired consistency and memory-less properties if multiple variables are sampled. Recursive sampling schemes have also been recently used to make progress on analyzing and designing fast sampling algorithms for a variety of Gibbs distributions (cf. [18, 7, 2]). As noted earlier, such schemes have been particularly useful for derandomising and constructing parallel and distributed versions of many popular MCMC algorithms for sampling solutions to CSPs [9, 11, 17, 18].
@@
We use these algorithms to carefully select and sample a small number of auxiliary variable assignments on an as-needed basis, showing that bad cases can be reduced to small calculations after a sublinear number of steps. Our proof of Section˜1 requires a novel adaptation of sampling techniques to avoid requiring global information or complete passes of the variable set; we show that we can perform analogous operations locally or leverage local uniformity properties to circumvent them altogether in the course of locally sampling a variable assignment. Importantly, we demonstrate how to execute local sampling in an oblivious, consistent fashion, so that the algorithm need not retain any memory and that variables need not be sampled in any particular order.
@@ 2 Preliminaries
Throughout, let $\Phi=(V,\mathcal{C})$ be a $k$-CNF on variable set $V=\{v_{1},\ldots,v_{n}\}$ with associated collection of clauses $\mathcal{C}$. In this work we do not let the same variable appear multiple times in any clause of $\Phi$, although our algorithm could be adapted to the more general scenario. We further assume that $\Phi$ is a $(k,d)$-formula; that is, each variable $v_{i}$ appears in at most $d$ clauses. For every clause $C\in\mathcal{C}$, let $\operatorname{vbl}(C)$ denote the collection of variables in the clause $C$. We further define $\Delta:=\max_{C\in\mathcal{C}}|\{C^{\prime}\in\mathcal{C}:\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset\}|$, so in particular $\Delta\leq kd$.
@@
[…]
Given a $(k,d)$-formula $\Phi$, let $H_{\Phi}=(V,\mathcal{E})$ be the dependency hypergraph (with multiple edges allowed), where $V$ is the set of variables and $\mathcal{E}=\{\operatorname{vbl}(C):C\in\mathcal{C}\}$ is the collection of variable sets of clauses of $\Phi$.
[…]
@@
Let $H_{\Phi}^{\sigma}$ be the associated (not necessarily $k$-uniform) hypergraph to $\Phi^{\sigma}$. For variable $v\in V\setminus\Lambda$, let $\Phi_{v}^{\sigma}$ denote the maximal connected component of $\Phi^{\sigma}$ to which $v$ belongs.
[…]
@@ 2.1 The random local access model and local computation algorithms
In this work, we construct a query-oblivious LCA for an intermediate step in our construction of the random local access model (described in more detail in Section˜2.2). We thus precisely define both LCAs and random local access models below.
[…]
@@
  • For a given input $\Pi$, upon being queried with $F_{i}$, the oracle with probability 1o(1)1-o(1) uses at most $t(n)$ resources (where $t(n)$ is sublinear in $n$) to return the value $\mathcal{A}(\Pi,\mathbf{R},F_{i}(Y))$ for some specific $Y\in\Omega$.
@@ 2.2 Marking and a query-oblivious LCA
As noted in the introduction, the Lovász local lemma [31] guarantees the existence of a satisfying assignment to any $(k,d)$-formula $\Phi$ if $d\leq(2ek)^{-1}2^{k}.$ Furthermore, the Moser-Tardos algorithm [28] gives a simple linear-time algorithm to construct such a satisfying assignment in the above $(k,d)$-regime:
  • Sample $v_{1},\ldots,v_{n}\overset{\textbf{R}}{\leftarrow}\{0,1\}$ uniformly at random;
@@
[…]
@@
[…]
@@
The construction of $\mathsf{IsMarked}(\cdot)$ and the verification that it is a query-oblivious LCA draws inspiration from the approach in [1] to obtain an oblivious LCA for hypergraph $2$-coloring. From a high level, $\mathsf{IsMarked}(\cdot)$ determines $s_{v}$ by randomly and greedily including variables in the marking and subsequently determining those that must/must not be in the marking, and finally (if needed) performing a rejection sampling on a smaller connected component that contains $v$. The formulae in Theorem˜2.1 are some technical conditions that guarantee this process to go through. We defer the proof of Theorem˜2.1 to Appendix˜E.
@@ 3 A random local access algorithm for $k$-SAT solutions
In this section, we introduce the main algorithm (Algorithm˜1) that locally accesses the assignment of a variable $v\in V$ in a uniformly random satisfying assignment $\sigma$ to $\Phi$.
Recall from Section˜1 that given a $(k,d)$-formula $\Phi=(V,\mathcal{C})$, variable $v\in V$, and any constant $c>0$, we wish to come up with a random local access algorithm $\mathcal{A}$ such that (1) the expected running time of $\mathcal{A}$ is $\mathrm{poly}\log(n)$, and (2) the output $\widehat{\mu}_{v}$ of $\mathcal{A}$ for every $v\in V$ satisfies $d_{\operatorname{\mathsf{TV}}}(\textsf{Law}(\widehat{\mu}_{v}),\mu_{v})\leq\frac{1}{n^{c}}$. As a high level description, given the above inputs, Algorithm˜1 performs the following:
@@
    Locally decides whether $v$ lies in an $\alpha$-marking of $\Phi$ using $\mathsf{IsMarked}(v)$ (Theorem˜2.1), such that the responses for all $v\in V$ yield a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
@@
    After obtaining the connected component $\Phi_{v}^{\sigma}$, we call $\mathsf{UniformSample}$ (Theorem˜3.2) to sample a uniformly random satisfying assignment to $\Phi_{v}^{\sigma}$ and extend $\sigma$. We then output $\sigma(v)$.
@@
To illustrate the workflow of Algorithm˜1, we present a toy example on a small $k$-CNF formula, omitting some subroutine details for brevity.
Example 3.1.
Suppose $k=3$, $d=2$ and $\alpha=1/3$. Consider the following $(k,d)$-formula $\Phi=(V,\mathcal{C})$ with variables $V=\{x_{1},x_{2},x_{3},x_{4},x_{5}\}$ and clauses $\mathcal{C}=\{C_{1},C_{2},C_{3}\}$, where
$C_{1}=(x_{1}\vee x_{2}\vee\neg x_{3}),\quad C_{2}=(\neg x_{2}\vee x_{3}\vee x_{4}),\quad C_{3}=(\neg x_{4}\vee x_{5}\vee\neg x_{1}).$
Suppose we wish to approximately sample $\sigma(x_{1})$ using the local access algorithm $\mathcal{A}$. We run $\mathsf{IsMarked}$ on each variable; suppose the resulting marking is $\mathcal{M}=\{x_{2},x_{5}\}$, so that $x_{1}$ is not marked.
Since $x_{1}\notin\mathcal{M}$, we call $\mathsf{Conn}(x_{1})$ to explore the connected component of $\Phi^{\mathcal{M}}$ containing $x_{1}$. We initialize $\mathcal{S}:=\{C_{1},C_{3}\}$ and partial assignment $\sigma=\emptyset$.
  • Process clause $C_{1}$. Since $x_{2}\in\mathcal{M}$, we call $\mathsf{MarginSample}(x_{2})$, which returns (say) $\sigma(x_{2})=0$. The clause becomes $(x_{1}\vee 0\vee\neg x_{3})=(x_{1}\vee\neg x_{3})$, which is unsatisfied with the current $\sigma$. We add adjacent clause $C_{2}$ (via $x_{2}$ and $x_{3}$) to $\mathcal{S}$.
  • Process clause $C_{3}$. Since $x_{5}\in\mathcal{M}$, we call $\mathsf{MarginSample}(x_{5})$, which returns (say) $\sigma(x_{5})=1$. The clause becomes $(\neg x_{4}\vee 1\vee\neg x_{1})$, which is satisfied.
  • Process clause $C_{2}$. We already have $\sigma(x_{2})=0$, so $\neg x_{2}=1$, and the clause is satisfied. Remove $C_{2}$ from $\mathcal{S}$.
Now the discovered component $\Phi_{x_{1}}^{\sigma}$ includes (1) clauses $C_{1}$, (2) marked variable assignments $\sigma(x_{2})=0$, $\sigma(x_{5})=1$, (3) free variables: $x_{1},x_{3},x_{4}$. We now run $\mathsf{UniformSample}$ on the reduced subformula $C_{1}^{\prime}=(x_{1}\vee\neg x_{3})$. The satisfying assignments $(x_{1},x_{3})$ are $\{(1,0,),(1,1),(0,0)\}$. We pick one uniformly at random, say $(x_{1},x_{3})=(1,0)$. Then we return $\widehat{\mu}_{x_{1}}=1$.
Algorithm˜1 is our main routine; it has several subroutines $\mathsf{IsMarked}$, $\mathsf{MarginSample}$, $\mathsf{Conn}$, and $\mathsf{UniformSample}$ that we now describe. Recall that $\mathsf{IsMarked}$ has been discussed in Theorem˜2.1. We now introduce the algorithm $\mathsf{UniformSample}$ given by the work of He–Wang–Yin [18].
@@
As seen in Theorem˜3.2, for an $n$-variable $(k,d)$-CNF formula, the algorithm has running time $O(n)$. However, as we will only apply $\mathsf{UniformSample}$ to connected components of size $\mathrm{poly}\log(n)$ (that are shattered by the partial assignment on the marked variables), the execution of $\mathsf{UniformSample}$ in Line 6 of Algorithm˜1 will only take polylogarithmic time.
We will define subroutines $\mathsf{MarginSample}$ and $\mathsf{Conn}$ below and show that they satisfy the desired correctness and efficiency properties, beginning by verifying $\mathsf{MarginSample}$ has the desired properties in Section˜4.
@@
Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking as in Section˜2.2. Fix positive constant $c>0$. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$, $36ed^{3}k^{4}\cdot 0.6^{\alpha k}\leq 1/2$, and $2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{\alpha k}}}\leq 0.9$. Then there exists a random local access algorithm $\mathsf{MarginSample}(\cdot)$ such that for every $u\in\mathcal{M}$, $\mathsf{MarginSample}(u)$ returns a random value in $\{0,1\}$ that satisfies the following:
@@
We also require $\mathsf{Conn}$ to be correct and have low expected cost.
@@
Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking as in Section˜2.2. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$ and $d\leq 2^{\alpha k/4}$. Then there exists a random local access algorithm $\mathsf{Conn}(\cdot)$ such that for every $v\in V\setminus\mathcal{M}$, $\mathsf{Conn}(v)$ returns the connected component in $\Phi^{\sigma}$ that contains $v$, where $\sigma$ is the partial assignment given by $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$. Moreover, for every $v\in V\setminus\mathcal{M}$, the expected cost of $\mathsf{Conn}(v)$ is $\mathrm{poly}\log(n)$.
From a high level, the algorithm $\mathsf{Conn}(v)$ explores the clauses and marked variables in the CNF formula that are reachable from $v$, greedily sampling the marked variables and expanding through unsatisfied clauses. It iteratively builds a partial assignment $\sigma$ that “shatters” the formula into disconnected components, isolating the one containing $v$. We will verify Theorem˜3.4 in Appendix˜D.
4 Proof of Theorem˜3.3
In this section we show Theorem˜3.3. Throughout, let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking with $|\mathcal{M}|=m$. We introduce a local access algorithm $\mathsf{MarginSample}$ that satisfies the key property that the joint distribution of outputs $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ consistently follows the marginal distribution $\mu_{\mathcal{M}}$. In particular, $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ will simulate the output of a systematic scan Glauber dynamics on the marked variables.
Definition 1.
Let $\mathcal{M}=\{u_{1},\ldots,u_{m}\}$ denote the marked variables (so $|\mathcal{M}|=m$).
@@
Input: a $k$-CNF formula $\Phi=(V,\mathcal{C})$, a set of marked variables $\mathcal{M}\subseteq V$, time $T$, and an ordering $\mathcal{M}=\{u_{1},\ldots,u_{m}\}$
@@
1:Sample $X_{0}(u)\overset{\textbf{R}}{\leftarrow}\{0,1\}$ uniformly and independently for each $u\in\mathcal{M}$
@@
We refer the readers to Appendix˜A for more introduction on the systematic scan Glauber dynamics and its comparison with the (original) projected Glauber dynamics Markov chain for sampling $k$-$\mathsf{SAT}$ solutions.